OpenStudy (anonymous):

8x + 13y = 2 y = x- 2 Solve the system of equations (4/3, -2/3) (4/21, -38/21) no solution correct?

OpenStudy (gorv):

y=x-2 put it in first equation 8x+13(x-2)=2 8x+13x-26=2 21x=28 x=28/21=4/3

OpenStudy (gorv):

y=4/3 - 2 = -2/3

OpenStudy (anonymous):

thanks